Multiple Versus Single Overhead Rates, Activity Drivers

Deoro Company has identified the following overhead activities, costs, and activity drivers for the coming year:

Activity Expected Cost Activity Driver Activity Capacity
Setting up equipment $468,460 Number of setups 590
Ordering costs 333,000 Number of orders 18,500
Machine costs 908,600 Machine hours 41,300
Receiving 336,000 Receiving hours 9,600

Deoro produces two models of dishwashers with the following expected prime costs and activity demands:

Model A Model B
Direct materials $577,000 $835,000
Direct labor $479,000 $480,000
Units completed 15,600 8,800
Direct labor hours 5,200 2,400
Number of setups 350 240
Number of orders 6,100 12,400
Machine hours 23,200 18,100
Receiving hours 3,600 6,000

1. Determine the unit cost for each model using direct labor hours to apply averhead. Round intermediate calculations and final answers to nearest cent.

Unit Cost

Model A ?

Model B ?

2. Determine the unit cost for each model using the four activity drivers. Round your answers to nearest cent.

Unit Cost

Model A ?

Model B ?

3. Which method produces the more accurate cost assignment?
Answer #1
Required 1 :
Total estimated overhead cost = 468460 + 333000 + 908600 + 336000 = 2046060
Predetermined overhead rate = Total estimated overhead cost / Total estimated direct labor hours = 2046060 / (5200+2400) = 269.22
Model A Model B
Direct materials 577000 835000
Direct labor 479000 480000
Overhead applied [ Predetermined overhead rate * Direct labor hours ] 1399944 646128
Total cost 2455944 1961128
(/) Units completed 15600 8800
Unit cost 157.43 222.86
Required 2 :
Activity Expected cost Expected activity Activity rate
Setting up equipment 468460 590 794
ordering cost 333000 18500 18
Machine costs 908600 41300 22
Receiving 336000 9600 35
Model A Model B
Direct materials 577000 835000
Direct labor 479000 480000
Overhead applied [ Activity rate * Activity ]
   Setting up equipment 277900 190560
   Ordering cost 109800 223200
   Machine costs 510400 398200
   Receiving 126000 210000
Total cost 2080100 2336960
(/) Units completed 15600 8800
Unit cost 133.34 265.56
3.
Which method produces the most accurate cost assigment?
Answer : Four activity drivers
Reason : As in four activity drivers the overhead amount is allocated in a better way as per the use of activity